40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Feb 14 17:40:27 2020, anchal, DailyProgress, ISS, Installed ISS on both paths using SR560s IMG_20200214_174127.jpgISS_Effect.pdf
    Reply  Sat Feb 15 18:57:43 2020, rana, DailyProgress, ISS, Installed ISS on both paths using SR560s 
       Reply  Thu Feb 20 17:42:23 2020, rana, DailyProgress, ISS, Installed New ISS on both paths using SR560s ISS_New_Analysis.pdf
          Reply  Sun Feb 23 18:19:32 2020, rana, DailyProgress, ISS, Installed New ISS on both paths using SR560s 
             Reply  Fri Feb 28 14:00:11 2020, anchal, DailyProgress, ISS, Installed New ISS on both paths using SR560s - PROBLEM 
                Reply  Sun Mar 1 16:34:31 2020, rana, DailyProgress, ISS, Installed New ISS on both paths using SR560s - PROBLEM 
                   Reply  Mon Mar 2 16:48:12 2020, anchal, DailyProgress, ISS, Follow-up, porbalem stopped happening, OLTF plot. NISS_OLTF.pdf
Message ID: 2545     Entry time: Fri Feb 28 14:00:11 2020     In reply to: 2544     Reply to this: 2546
Author: anchal 
Type: DailyProgress 
Category: ISS 
Subject: Installed New ISS on both paths using SR560s - PROBLEM 

I'm currently struggling with a problem in North ISS which I think needs to be documented here. Here's the synopsis:


The problem:

  • After running for the weekend with high gains and gain peaking at ~45 kHz, I found this week that the North ISS was overloading.
  • The power level reaching the cavities change over the course of the day by +- 20-30% usually (still a mystery to me), so I assumed this could just be because of increased laser power saturating the SR560.
  • But what I have found is that except for some random instances, I can't keep the gain high enough in this ISS path.
  • At the start, I thought this problem is at least reset-able by unlocking the North FSS and locking the cavity back. But later, I've found that it is not working either now.
  • The defining feature is, as I slowly increase the gain (giving ample time for the loop to settle), at the first point where I see any sign of overload, it happens at a rate of ~300 mHz, that's once every 3  seconds!
  • I checked at the photodiode output at this gain setting and I see the loop gets unstable and for about half a second every 3 seconds or so.
  • On increasing the gain further, this overload just becomes a constant condition.
  • My first guess is that somehow the lower unity gain frequency is becoming unstable. But the mystery is:
    • Why did this not happen before?
    • Why is this not seen in the South Path at all? It is healthy as ever.
    • What really changed during the weekend?
    • And why once in a while, it was actually working?
  • The only difference in the SR560 settings for the two loops is that the South path has 'invert' ON while North path does not. At first, I was perplexed by this too. South needed 'invert' to be stable and the North did not. Maybe this is the reason?
  • I have ruled out any failure in SR560 by switching it with another one and seeing the exact same phenomenon.
  • Remaining suspects are a damaged BNC cable/connector, damaged photodiode or EOAM (I wish not). However, some bad control loop configuration is my primary suspect.

I would like it if anyone has any comments on any of the points above or suggestions to tackle this problem. I can make some measurements and post them on requests.

ELOG V3.1.3-